Grade 12General Physics

A body of mass m is dropped feom a height h on a sand floor .If the body pentrates x cm on the sand the average resistance offered by the sand to the body is

janvi
8 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Wheb body is dropped on a sand floor
mgh = ½ mv²
Now if F is testing force then
Resultant force = F - mg
Hence
(F -mg)x = mgh
 Hence
F = mg (1 + h/x)
Regards
Arun (askIITiams forum expert)